# PINSAT: Parallelized Interleaving of Graph Search and Trajectory Optimization for Kinodynamic Motion Planning

### 17 Jan 2024 v1

@@ 6 Evaluation
Table 1 shows success rate and planning time statistics for INSAT and PINSAT for different thread budgets. The planning time statistics were only computed over the problems successfully solved by both algorithms for a given thread budget. We also implemented a parallelized version of RRTConnect (pBiRRT) as a baseline. The kinematic plan generated by pBiRRT was then post-processed by the same B-spline optimization used in PINSAT to compute the final kinodynamic plan. As the optimization has a constraint on the maximum execution duration of the trajectory, adding all the waypoints from the path at once will result in an over-constrained system of equations for which a solution may not exist. To circumvent this, we iteratively add the waypoint constraint starting from the minimum containing just the start and the goal state. Though pBiRRT had a 100% success rate in computing the spatial trajectory, the results after post-processing with the optimizer were abysmal. PINSAT achieves a significantly higher success rate than INSAT for all thread budgets greater than 1. Specifically, it achieves a 5x improvement in mean planning time, a 7x improvement in median planning time, and a 1.8x improvement in success rate for $N_{t}$ = 60. Even with a single thread, PINSAT achieves lower planning times than INSAT. This is because of the decoupling of edge evaluations from state expansions in ePA*SE.
7 Conclusion
We presented PINSAT, which parallelizes the interleaving of search and trajectory optimization using the ideas of ePA*SE (Mukherjee, Aine, and Likhachev 2022a). In a kinodynamic manipulation planning domain, PINSAT achieved significantly higher success rates than INSAT (Natarajan, Choset, and Likhachev 2021; Natarajan et al. 2023) and a drastic reduction in planning time.

### 16 Mar 2024 v2

@@ VI Evaluation
Table I shows success rate and planning time statistics for INSAT and PINSAT for different thread budgets. The planning time statistics were only computed over the problems successfully solved by both algorithms for a given thread budget. We also implemented a parallelized version of RRTConnect (pBiRRT) as a baseline. The kinematic plan generated by pBiRRT was then post-processed by the same B-spline optimization used in PINSAT to compute the final kinodynamic plan. As the optimization has a constraint on the maximum execution duration of the trajectory, adding all the waypoints from the path at once will result in an over-constrained system of equations for which a solution may not exist. To circumvent this, we iteratively add the waypoint constraint starting from the minimum containing just the start and the goal state. Though pBiRRT had a 100% success rate in computing the spatial trajectory, the results after post-processing with the optimizer were abysmal. PINSAT achieves a significantly higher success rate than INSAT for all thread budgets greater than 1. Specifically, it achieves a 5x improvement in mean planning time, a 7x improvement in median planning time, and a 1.8x improvement in success rate for $N_{t}$ = 50. Even with a single thread, PINSAT achieves lower planning times than INSAT. This is because of the decoupling of edge evaluations from state expansions in ePA*SE.
VII Conclusion
We presented PINSAT, an extension of INSAT that asynchronously parallelizes the numerous expensive calls to the trajectory optimizer. PINSAT achieves this by incorporating the concepts of edge expansion and asynchronous edge evaluation borrowed from the recently introduced parallel search algorithm ePA*SE [8]. We evaluated our algorithm using a kinodynamic manipulation planning domain and demonstrated significantly higher success rates than INSAT [6, 7], along with a drastic reduction in planning time. PINSAT guarantees only completeness at present, but in the future, we aim to explore how to establish independence checks, as done in ePA*SE, to guarantee optimality.”